OpenStudy (anonymous):

WILL MEDAL AND FAN YOU !!!!! 1/4(3x-16)+1/4x=5 1/3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do u want it simplified?

OpenStudy (anonymous):

i need a whole number and a fraction answer

OpenStudy (misty1212):

hi!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but theirs no whole number

OpenStudy (anonymous):

theres*

OpenStudy (misty1212):

it would be easier if you multiplied both sides by 4 and started with \[3x-16+x=4\times 5\tfrac{ 2}{3}\]

OpenStudy (anonymous):

ok 17/3 how many times does 3 go into 17

OpenStudy (anonymous):

5

OpenStudy (anonymous):

close 5.67 times

OpenStudy (anonymous):

so that's your answer

OpenStudy (anonymous):

but it needs to be a fraction. so if i put 5.67 into a fraction the fraction would be 5 67⁄100 that would be my answer ?

OpenStudy (anonymous):

my bad 5 67/100

OpenStudy (anonymous):

yes then u got to simplify it

OpenStudy (anonymous):

14/25?

OpenStudy (anonymous):

\[5\frac{ 67 }{ 100 }=5\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

wait thats not right

OpenStudy (anonymous):

ok so i redid it and i think the answer is 9 1/3

OpenStudy (anonymous):

no its 5 something

OpenStudy (anonymous):

i dont think it is.

OpenStudy (anonymous):

67/100 cant be simplified

OpenStudy (anonymous):

lemme summit 9 1/3 and see

OpenStudy (anonymous):

i was right (: 9 1/3 is the answer

OpenStudy (anonymous):

thats weird ok sorry

OpenStudy (anonymous):

its fine still medal'd & fanned you for trying to help (:

OpenStudy (anonymous):

ok thanks